Patients arrive at hospital accident and emergency at random at an average rate of 6 per hour. Find the probability that, during any 60-minute period, the number of patients arriving at the hospital accident and emergency department is at least 2. Round off answers to 4 decimal places.




1
Expert's answer
2021-02-24T06:05:26-0500

This is a Poisson distribution with "\lambda=6."

"P(X\ge2)=1-P(X=0)-P(X=1)=1-e^{-6}(\frac{6^0}{0!}+\frac{6^1}{1!})=0.9827."